n!(n+2)=n!+(n+1)!

Answer by LinnW(1048)

n!(n+2)=n!+(n+1)!
If we rewrite n+2 to (n+1) + 1
we get
n!( (n+1) + 1 ) = n! + (n+1)!
Since (n+1)n! = (n+1)!
(n+1)! + n! = n! + (n+1)!
So it appears that for any integer n the equality holds.